# Towards Principled Assessment of Tabular Data Synthesis Algorithms

### v1

@@ Appendix A Hyperparameters Search Spaces of synthesis Algorithms
Table A1: PGM [41] hyperparameters search space.
@@
Table A2: PrivSyn [70] hyperparameters search space.
@@
Table A3: TVAE [65] hyperparameters search space.
@@
Table A4: CTGAN [65] hyperparameters search space.
@@
Table A5: TabDDPM [32] hyperparameters search space.
@@
Table A6: GReaT [7] hyperparameters search space.
@@
Table A7: PATE-GAN [30] hyperparameters search space.
@@
Table A8: TableDiffusion [60] hyperparameters search space.
@@ Appendix C Hyperparameters Search Spaces of Machine Learning Models
Table C9: SVM hyperparameters search space.
@@
Table C10: Logistic regression (Ridge regression) hyperparameters search space.
@@
Table C11: Decision tree hyperparameters search space.
@@
Table C12: Random forest hyperparameters search space.
@@
Table C13: MLP hyperparameters search space.
@@
Table C14: XGBoost hyperparameters search space.
@@
Table C15: CatBoost hyperparameters search space from [22].
@@
Table C16: Transformer hyperparameters space from [22].
@@
Appendix D Complete Experimental Results
D.1 Complete Fidelity Results
Here we include the remaining fidelity results in our evaluation. The overall fidelity evaluation on test data is demonstrated in Table D17.
@@ D.1 Complete Fidelity Results
For fidelity on test data $\mathcal{D}_{\text{train}}$, the detailed Wasserstein distances for one-way marginals are shown in Table D18 and Table D19. The Wasserstein distances for two-way marginals are shown in Table D20, Table D21 and Table D22.
For fidelity on test data $\mathcal{D}_{\text{test}}$, the detailed Wasserstein distances for one-way marginals are shown in Table D23 and Table D24. The Wasserstein distances for two-way marginals are shown in Table D25, Table D26 and Table D27.
Table D17: Fidelity evaluation of data synthesis on test data $\mathcal{D}_{\text{test}}$. The privacy budget $\epsilon$ of HP synthesis algorithms is set as $\infty$ (the top part), and the budget for DP synthesis algorithm is set as $1$ (the bottom part). The best result of each category is in bold.
@@
Table D18: Fidelity evaluation of one-way categorical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
@@
Table D19: Fidelity evaluation of one-way numerical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
@@
Table D20: Fidelity evaluation of two-way categorical-categorical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
@@
Table D21: Fidelity evaluation of two-way categorical-numerical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
@@
Table D22: Fidelity evaluation of two-way numerical-numerical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
@@
Table D23: Fidelity evaluation of one-way categorical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
@@
Table D24: Fidelity evaluation of one-way numerical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
@@
Table D25: Fidelity evaluation of two-way categorical-categorical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
@@
Table D26: Fidelity evaluation of two-way categorical-numerical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
@@
Table D27: Fidelity evaluation of two-way numerical-numerical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
@@
D.2 Utility Results
Here we show the utility results for all synthesizers. Specifically, the machine learning affinity results are shown in Table D28, and the query results are detailed in Table D29.
Table D28: Machine learning affinity of data synthesis. The privacy budget $\epsilon$ of HP synthesizers is set as $\infty$ (the top part), and the budget for DP synthesizers is set as $1$ (the bottom part). The lower value means the performance of synthetic data is more similar to real data. The best result of each category is in bold.
[…]
Table D29: Query error of data synthesis. The privacy budget $\epsilon$ of HP synthesizers is set as $\infty$ (the top part), and the budget for DP synthesizers is set as $1$ (the bottom part). The best result of each category is in bold.
Appendix E Misleading conclusion

### v2

@@ Appendix A Hyperparameters Search Spaces of synthesis Algorithms
Table A1. PGM (McKenna et al., 2019) hyperparameters search space.
@@
Table A2. PrivSyn (Zhang et al., 2021) hyperparameters search space.
@@
Table A3. TVAE (Xu et al., 2019) hyperparameters search space.
@@
Table A4. CTGAN (Xu et al., 2019) hyperparameters search space.
@@
Table A5. TabDDPM (Kotelnikov et al., 2023) hyperparameters search space.
@@
Table A6. GReaT (Borisov et al., 2023) hyperparameters search space.
@@
Table A7. PATE-GAN (Jordon et al., 2018) hyperparameters search space.
@@
Table A8. TableDiffusion (Truda, 2023) hyperparameters search space.
@@ Appendix C Hyperparameters Search Spaces of Machine Learning Models
Table C9. SVM hyperparameters search space.
@@
Table C10. Logistic regression (Ridge regression) hyperparameters search space.
@@
Table C11. Decision tree hyperparameters search space.
@@
Table C12. Random forest hyperparameters search space.
@@
Table C13. MLP hyperparameters search space.
@@
Table C14. XGBoost hyperparameters search space.
@@
Table C15. CatBoost hyperparameters search space from (Gorishniy et al., 2021).
@@
Table C16. Transformer hyperparameters space from (Gorishniy et al., 2021).
@@
Appendix D Implementation of existing privacy metrics
We implement two existing private metrics (i.e., DCR (Zhao et al., 2021) and MI attack (Stadler et al., 2022)) for performance comparison. For DCR, we first normalize all attributes of synthetic data and real data, and calculate the distance distribution from each synthetic data point to its nearest real one, using the 5th percentile as the privacy score. For the MI attack, we follow (Stadler et al., 2022) to conduct feature extraction from synthetic data for shadow model training. We use the most widely-used metric: true-positive rate (TPR) at 10% false-positive rates (FPR) (Carlini et al., 2022), i.e., , TPR@10% as the attack score for comparison.
Appendix E Complete Experimental Results
E.1. Complete Fidelity Results
Here we include the remaining fidelity results in our evaluation. The overall fidelity evaluation on test data is demonstrated in Table E17.
@@ E.1. Complete Fidelity Results
For fidelity on test data $\mathcal{D}_{\text{train}}$, the detailed Wasserstein distances for one-way marginals are shown in Table E18 and Table E19. The Wasserstein distances for two-way marginals are shown in Table E20, Table E21 and Table E22.
For fidelity on test data $\mathcal{D}_{\text{test}}$, the detailed Wasserstein distances for one-way marginals are shown in Table E23 and Table E24. The Wasserstein distances for two-way marginals are shown in Table E25, Table E26 and Table E27.
Table E17. Fidelity evaluation (i.e., Wasserstein distance) of data synthesis algorithms on test data $\mathcal{D}_{\text{test}}$. The privacy budget $\epsilon$ of HP synthesizers is $\infty$ (the top part), and the budget for DP synthesizers is $1$ (the middle part). HALF and HISTOGRAM are the baselines that serve as the empirical upper bound and lower bound of the fidelity for HP synthesizers. The low score indicates the synthesizer can generate high-quality synthetic data that preserves the marginals of real ones. The best result is in bold.
@@
HALF $0.022\scriptscriptstyle\pm\scriptstyle.002$ $0.023\scriptscriptstyle\pm\scriptstyle.002$ $0.016\scriptscriptstyle\pm\scriptstyle.003$ $0.011\scriptscriptstyle\pm\scriptstyle.003$ $0.042\scriptscriptstyle\pm\scriptstyle.005$ $0.015\scriptscriptstyle\pm\scriptstyle.003$ $0.041\scriptscriptstyle\pm\scriptstyle.006$ $0.023\scriptscriptstyle\pm\scriptstyle.005$ $0.028\scriptscriptstyle\pm\scriptstyle.007$ $0.010\scriptscriptstyle\pm\scriptstyle.002$ $0.060\scriptscriptstyle\pm\scriptstyle.007$ $0.014\scriptscriptstyle\pm\scriptstyle.001$
HISTOGRAM $0.199\scriptscriptstyle\pm\scriptstyle.017$ $0.1$001 $0.030\scriptscriptstyle\pm\scriptstyle.001$ $0.048\scriptscriptstyle\pm\scriptstyle.002$ $0.113\scriptscriptstyle\pm\scriptstyle.006$ $0.080\scriptscriptstyle\pm\scriptstyle.003$ $0.066\scriptscriptstyle\pm\scriptstyle.002$ $0.065\scriptscriptstyle\pm\scriptstyle.002$ $0.094\scriptscriptstyle\pm\scriptstyle.009$ $0.059\scriptscriptstyle\pm\scriptstyle.006$ $0.081\scriptscriptstyle\pm\scriptstyle.004$ $0.076\scriptscriptstyle\pm\scriptstyle.001$
Table E18. Fidelity evaluation of one-way categorical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
@@
HALF $0.019\scriptscriptstyle\pm\scriptstyle.002$ $0.018\scriptscriptstyle\pm\scriptstyle.001$ $0.008\scriptscriptstyle\pm\scriptstyle.002$ $0.010\scriptscriptstyle\pm\scriptstyle.007$ $0.014\scriptscriptstyle\pm\scriptstyle.002$ $0.022\scriptscriptstyle\pm\scriptstyle.006$ $0.012\scriptscriptstyle\pm\scriptstyle.003$ $0.008\scriptscriptstyle\pm\scriptstyle.002$ $0.015\scriptscriptstyle\pm\scriptstyle.006$ $0.006\scriptscriptstyle\pm\scriptstyle.001$ $0.021\scriptscriptstyle\pm\scriptstyle.004$
HISTOGRAM $0.017\scriptscriptstyle\pm\scriptstyle.002$ $0.015\scriptscriptstyle\pm\scriptstyle.001$ $0.008\scriptscriptstyle\pm\scriptstyle.001$ $0.010\scriptscriptstyle\pm\scriptstyle.003$ $0.008\scriptscriptstyle\pm\scriptstyle.002$ $0.032\scriptscriptstyle\pm\scriptstyle.004$ $0.008\scriptscriptstyle\pm\scriptstyle.001$ $0.006\scriptscriptstyle\pm\scriptstyle.004$ $0.008\scriptscriptstyle\pm\scriptstyle.002$ $0.006\scriptscriptstyle\pm\scriptstyle.001$ $0.016\scriptscriptstyle\pm\scriptstyle.001$
Table E19. Fidelity evaluation of one-way numerical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
@@
HALF $0.004\scriptscriptstyle\pm\scriptstyle.000$ $0.003\scriptscriptstyle\pm\scriptstyle.001$ $0.004\scriptscriptstyle\pm\scriptstyle.001$ $0.006\scriptscriptstyle\pm\scriptstyle.001$ $0.004\scriptscriptstyle\pm\scriptstyle.001$ $0.008\scriptscriptstyle\pm\scriptstyle.001$ $0.005\scriptscriptstyle\pm\scriptstyle.000$ $0.004\scriptscriptstyle\pm\scriptstyle.001$ $0.004\scriptscriptstyle\pm\scriptstyle.000$ $0.009\scriptscriptstyle\pm\scriptstyle.002$ $0.003\scriptscriptstyle\pm\scriptstyle.000$
HISTOGRAM $0.226\scriptscriptstyle\pm\scriptstyle.016$ $0.099\scriptscriptstyle\pm\scriptstyle.003$ $0.035\scriptscriptstyle\pm\scriptstyle.002$ $0.037\scriptscriptstyle\pm\scriptstyle.000$ $0.035\scriptscriptstyle\pm\scriptstyle.001$ $0.039\scriptscriptstyle\pm\scriptstyle.000$ $0.041\scriptscriptstyle\pm\scriptstyle.000$ $0.034\scriptscriptstyle\pm\scriptstyle.000$ $0.068\scriptscriptstyle\pm\scriptstyle.002$ $0.029\scriptscriptstyle\pm\scriptstyle.001$ $0.042\scriptscriptstyle\pm\scriptstyle.000$
Table E20. Fidelity evaluation of two-way categorical-categorical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
@@
HALF $0.045\scriptscriptstyle\pm\scriptstyle.004$ $0.040\scriptscriptstyle\pm\scriptstyle.002$ $0.014\scriptscriptstyle\pm\scriptstyle.002$ $0.027\scriptscriptstyle\pm\scriptstyle.002$ $0.023\scriptscriptstyle\pm\scriptstyle.004$ $0.011\scriptscriptstyle\pm\scriptstyle.001$ $0.041\scriptscriptstyle\pm\scriptstyle.006$
HISTOGRAM $0.116\scriptscriptstyle\pm\scriptstyle.002$ $0.072\scriptscriptstyle\pm\scriptstyle.002$ $0.046\scriptscriptstyle\pm\scriptstyle.001$ $0.188\scriptscriptstyle\pm\scriptstyle.005$ $0.053\scriptscriptstyle\pm\scriptstyle.002$ $0.027\scriptscriptstyle\pm\scriptstyle.002$ $0.044\scriptscriptstyle\pm\scriptstyle.003$
Table E21. Fidelity evaluation of two-way categorical-numerical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
@@
HALF $0.025\scriptscriptstyle\pm\scriptstyle.002$ $0.022\scriptscriptstyle\pm\scriptstyle.001$ $0.015\scriptscriptstyle\pm\scriptstyle.007$ $0.021\scriptscriptstyle\pm\scriptstyle.002$ $0.028\scriptscriptstyle\pm\scriptstyle.006$ $0.021\scriptscriptstyle\pm\scriptstyle.004$ $0.016\scriptscriptstyle\pm\scriptstyle.001$ $0.020\scriptscriptstyle\pm\scriptstyle.005$ $0.011\scriptscriptstyle\pm\scriptstyle.001$ $0.033\scriptscriptstyle\pm\scriptstyle.005$
HISTOGRAM $0.253\scriptscriptstyle\pm\scriptstyle.015$ $0.117\scriptscriptstyle\pm\scriptstyle.003$ $0.060\scriptscriptstyle\pm\scriptstyle.003$ $0.078\scriptscriptstyle\pm\scriptstyle.001$ $0.163\scriptscriptstyle\pm\scriptstyle.002$ $0.064\scriptscriptstyle\pm\scriptstyle.001$ $0.085\scriptscriptstyle\pm\scriptstyle.003$ $0.094\scriptscriptstyle\pm\scriptstyle.001$ $0.076\scriptscriptstyle\pm\scriptstyle.002$ $0.061\scriptscriptstyle\pm\scriptstyle.001$
Table E22. Fidelity evaluation of two-way numerical-numerical marginals on train data $\mathcal{D}_{\text{train}}$. The best result of each category is in bold.
@@
HALF $0.009\scriptscriptstyle\pm\scriptstyle.001$ $0.007\scriptscriptstyle\pm\scriptstyle.002$ $0.013\scriptscriptstyle\pm\scriptstyle.001$ $0.014\scriptscriptstyle\pm\scriptstyle.001$ $0.012\scriptscriptstyle\pm\scriptstyle.001$ $0.021\scriptscriptstyle\pm\scriptstyle.002$ $0.013\scriptscriptstyle\pm\scriptstyle.001$ $0.009\scriptscriptstyle\pm\scriptstyle.002$ $0.011\scriptscriptstyle\pm\scriptstyle.001$ $0.025\scriptscriptstyle\pm\scriptstyle.003$ $0.008\scriptscriptstyle\pm\scriptstyle.000$
HISTOGRAM $0.453\scriptscriptstyle\pm\scriptstyle.031$ $0.205\scriptscriptstyle\pm\scriptstyle.005$ $0.098\scriptscriptstyle\pm\scriptstyle.004$ $0.096\scriptscriptstyle\pm\scriptstyle.001$ $0.120\scriptscriptstyle\pm\scriptstyle.002$ $0.093\scriptscriptstyle\pm\scriptstyle.001$ $0.113\scriptscriptstyle\pm\scriptstyle.000$ $0.142\scriptscriptstyle\pm\scriptstyle.000$ $0.140\scriptscriptstyle\pm\scriptstyle.003$ $0.080\scriptscriptstyle\pm\scriptstyle.002$ $0.094\scriptscriptstyle\pm\scriptstyle.000$
Table E23. Fidelity evaluation of one-way categorical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
@@
HALF $0.020\scriptscriptstyle\pm\scriptstyle.003$ $0.023\scriptscriptstyle\pm\scriptstyle.003$ $0.012\scriptscriptstyle\pm\scriptstyle.003$ $0.011\scriptscriptstyle\pm\scriptstyle.005$ $0.037\scriptscriptstyle\pm\scriptstyle.007$ $0.019\scriptscriptstyle\pm\scriptstyle.005$ $0.030\scriptscriptstyle\pm\scriptstyle.005$ $0.018\scriptscriptstyle\pm\scriptstyle.009$ $0.028\scriptscriptstyle\pm\scriptstyle.008$ $0.008\scriptscriptstyle\pm\scriptstyle.002$ $0.048\scriptscriptstyle\pm\scriptstyle.011$
HISTOGRAM $0.017\scriptscriptstyle\pm\scriptstyle.002$ $0.020\scriptscriptstyle\pm\scriptstyle.001$ $0.011\scriptscriptstyle\pm\scriptstyle.001$ $0.011\scriptscriptstyle\pm\scriptstyle.003$ $0.046\scriptscriptstyle\pm\scriptstyle.010$ $0.017\scriptscriptstyle\pm\scriptstyle.005$ $0.025\scriptscriptstyle\pm\scriptstyle.001$ $0.012\scriptscriptstyle\pm\scriptstyle.004$ $0.038\scriptscriptstyle\pm\scriptstyle.020$ $0.009\scriptscriptstyle\pm\scriptstyle.001$ $0.047\scriptscriptstyle\pm\scriptstyle.003$
Table E24. Fidelity evaluation of one-way numerical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
@@
HALF $0.005\scriptscriptstyle\pm\scriptstyle.001$ $0.003\scriptscriptstyle\pm\scriptstyle.000$ $0.004\scriptscriptstyle\pm\scriptstyle.001$ $0.015\scriptscriptstyle\pm\scriptstyle.002$ $0.004\scriptscriptstyle\pm\scriptstyle.001$ $0.018\scriptscriptstyle\pm\scriptstyle.003$ $0.011\scriptscriptstyle\pm\scriptstyle.001$ $0.013\scriptscriptstyle\pm\scriptstyle.003$ $0.005\scriptscriptstyle\pm\scriptstyle.001$ $0.018\scriptscriptstyle\pm\scriptstyle.003$ $0.007\scriptscriptstyle\pm\scriptstyle.001$
HISTOGRAM $0.208\scriptscriptstyle\pm\scriptstyle.020$ $0.094\scriptscriptstyle\pm\scriptstyle.001$ $0.033\scriptscriptstyle\pm\scriptstyle.000$ $0.051\scriptscriptstyle\pm\scriptstyle.001$ $0.034\scriptscriptstyle\pm\scriptstyle.000$ $0.043\scriptscriptstyle\pm\scriptstyle.001$ $0.042\scriptscriptstyle\pm\scriptstyle.000$ $0.047\scriptscriptstyle\pm\scriptstyle.002$ $0.061\scriptscriptstyle\pm\scriptstyle.007$ $0.039\scriptscriptstyle\pm\scriptstyle.004$ $0.048\scriptscriptstyle\pm\scriptstyle.001$
Table E25. Fidelity evaluation of two-way categorical-categorical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
@@
HALF $0.049\scriptscriptstyle\pm\scriptstyle.003$ $0.052\scriptscriptstyle\pm\scriptstyle.004$ $0.021\scriptscriptstyle\pm\scriptstyle.003$ $0.060\scriptscriptstyle\pm\scriptstyle.008$ $0.056\scriptscriptstyle\pm\scriptstyle.009$ $0.013\scriptscriptstyle\pm\scriptstyle.003$ $0.099\scriptscriptstyle\pm\scriptstyle.008$
HISTOGRAM $0.116\scriptscriptstyle\pm\scriptstyle.002$ $0.081\scriptscriptstyle\pm\scriptstyle.000$ $0.049\scriptscriptstyle\pm\scriptstyle.001$ $0.217\scriptscriptstyle\pm\scriptstyle.007$ $0.072\scriptscriptstyle\pm\scriptstyle.002$ $0.030\scriptscriptstyle\pm\scriptstyle.001$ $0.096\scriptscriptstyle\pm\scriptstyle.003$
Table E26. Fidelity evaluation of two-way categorical-numerical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
@@
HALF $0.027\scriptscriptstyle\pm\scriptstyle.003$ $0.029\scriptscriptstyle\pm\scriptstyle.003$ $0.016\scriptscriptstyle\pm\scriptstyle.005$ $0.056\scriptscriptstyle\pm\scriptstyle.006$ $0.025\scriptscriptstyle\pm\scriptstyle.006$ $0.052\scriptscriptstyle\pm\scriptstyle.008$ $0.035\scriptscriptstyle\pm\scriptstyle.008$ $0.043\scriptscriptstyle\pm\scriptstyle.010$ $0.013\scriptscriptstyle\pm\scriptstyle.002$ $0.077\scriptscriptstyle\pm\scriptstyle.011$
HISTOGRAM $0.236\scriptscriptstyle\pm\scriptstyle.019$ $0.117\scriptscriptstyle\pm\scriptstyle.001$ $0.053\scriptscriptstyle\pm\scriptstyle.003$ $0.123\scriptscriptstyle\pm\scriptstyle.009$ $0.151\scriptscriptstyle\pm\scriptstyle.005$ $0.084\scriptscriptstyle\pm\scriptstyle.001$ $0.092\scriptscriptstyle\pm\scriptstyle.003$ $0.122\scriptscriptstyle\pm\scriptstyle.012$ $0.071\scriptscriptstyle\pm\scriptstyle.007$ $0.106\scriptscriptstyle\pm\scriptstyle.005$
Table E27. Fidelity evaluation of two-way numerical-numerical marginals on test data $\mathcal{D}_{\text{test}}$. The best result of each category is in bold.
@@
HALF $0.011\scriptscriptstyle\pm\scriptstyle.001$ $0.008\scriptscriptstyle\pm\scriptstyle.000$ $0.013\scriptscriptstyle\pm\scriptstyle.001$ $0.039\scriptscriptstyle\pm\scriptstyle.003$ $0.013\scriptscriptstyle\pm\scriptstyle.001$ $0.050\scriptscriptstyle\pm\scriptstyle.004$ $0.028\scriptscriptstyle\pm\scriptstyle.001$ $0.027\scriptscriptstyle\pm\scriptstyle.005$ $0.012\scriptscriptstyle\pm\scriptstyle.001$ $0.059\scriptscriptstyle\pm\scriptstyle.004$ $0.020\scriptscriptstyle\pm\scriptstyle.001$
HISTOGRAM $0.417\scriptscriptstyle\pm\scriptstyle.041$ $0.194\scriptscriptstyle\pm\scriptstyle.003$ $0.093\scriptscriptstyle\pm\scriptstyle.001$ $0.127\scriptscriptstyle\pm\scriptstyle.003$ $0.118\scriptscriptstyle\pm\scriptstyle.001$ $0.104\scriptscriptstyle\pm\scriptstyle.002$ $0.116\scriptscriptstyle\pm\scriptstyle.000$ $0.168\scriptscriptstyle\pm\scriptstyle.003$ $0.125\scriptscriptstyle\pm\scriptstyle.015$ $0.115\scriptscriptstyle\pm\scriptstyle.006$ $0.104\scriptscriptstyle\pm\scriptstyle.002$
Appendix F Discussion of proposed privacy metric
The definition of the proposed membership inference score (MDS) aligns closely with concepts of memorization in neural networks (Feldman, 2020; Zhang et al., 2023a) and the leave-one-out notion of stability in machine learning (Bousquet and Elisseeff, 2002). However, it diverges in three crucial ways: (i) We measure the worst-case discourse risk as the privacy metric, whereas other studies focus on the difference of individuals or average cases. (ii) Our work specifically addresses privacy concerns in data synthesis, as opposed to other studies that explore discriminative models like classification. (iii) Our approach emphasizes the discrepancy caused by the presence or absence of target data in training, in contrast to other works that highlight performance gains from adding samples to the training set.
Appendix G Misleading conclusion